OpenStudy (anonymous):

Solve for y. y/9 = 4 2/3

OpenStudy (anonymous):

4 2/3 = 14/3 14/3 * 9 = y 14/3*9 = 126/3 = 42 y = 42

OpenStudy (anonymous):

thx :D